OpenStudy (rockycatsrule):

What is the solution to the system of equations represented by the two equations? y= 3/2x y= -1/2x+4

pooja195 (pooja195):

y= 3/2x <----Equation 1 y= -1/2x+4 <----Equation 2 We know the value of y already its given so take equation 1 and plug it into equation 2 \[\huge~\rm~3/2x=-1/2x+4 \] solve for x

OpenStudy (rockycatsrule):

so it would be (3,2)

OpenStudy (anonymous):

i think it might be (2, 3) instead @rockycatsrule

OpenStudy (rockycatsrule):

ok thx

OpenStudy (anonymous):

because when i solved for x it came out as 2